• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 00:04:19 on June 4,2004

    Entry number 124655

    keyword=Shift Summary

    Program is production running on kin 17
    Left arm negative P0=2.06 GeV/c
    Right arm negative P0= 2.00 GeV/c
    15 cm LH2

    18:09 DAQ problem had to reboot rocs 12 and 16, kcoda and restart (16 minutes)
    19:24 Indication of ROC problem at end of run 1697, but runstart worked
    ~22:00 Bogdan brings us pizza! Thanks Bogdan!
    Very good beam!
    Smooth runninig!

    Run summary
    1686 16:00-16:20 1.0M
    1687 16:20-16:43 1.0M
    1688 16:43-17:04 1.0M
    1689 17:05-17:24 1.0M
    1690 17:24-17:48 1.0M
    1691 17:49-18:09 1.0M
    had to reboot ROC12 and 16 after end run. Run numbers 1692 and 1693 skipped
    1694 18:25-18:43 1.0M
    1695 18:43-19:04 1.0M
    1696 19:04-19:24 1.0M Indication of ROC problem at end of run, but runstart worked
    1697 19:25-19:46 1.0M
    1698 19:46-20:07 1.0M
    1699 20:07-20:29 1.0M
    1700 20:29-20:49 1.0M
    1701 20:50-21:11 1.0M
    1702 21:12-21:31 1.0M
    1703 21:32-21:54 1.0M
    1704 21:55-22:14 1.0M
    1705 22:15-22:37 1.0M
    1706 22:37-22:57 1.0M
    1707 22:57-23:20 1.0M
    1708 23:20-23:46 1.0M (21M)
    1709 23:46 continues into owl shift